OpenStudy (anonymous):

Help! :) Insert three arithmetic means between 11 and 14.

OpenStudy (blockcolder):

11 (1) (2) (3) 14 (1)=11+d (2)=11+2d (3)=11+3d 14=11+4d Solve this last equation for d, and plug them in (1), (2), and (3).

OpenStudy (anonymous):

d = 3/4 ah.. got it :) thanks! ^_^

OpenStudy (blockcolder):

No problem. :)

OpenStudy (anonymous):

Eh how about this po, On five separate examinations, a student received the ff. test scores: 78, 89, 62, 75 and 84. Find the arithmetic mean of these scores. How to solve this?

OpenStudy (blockcolder):

Arithmetic mean is simply the average of all the items. Add them all, then divide by how many items there are.

OpenStudy (anonymous):

ah.. okay. :) thanks again :)